John can clean the pool in 30 minutes, Julie can clean the pool in 45 minutes and Tom can clean the pool in 90 minutes. How long would it take to clean the pool if all 3 work together
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

John can clean the pool in 30 minutes, Julie can clean the pool in 45 minutes and Tom can clean the pool in 90 minutes. How long would it take to clean the pool if all 3 work together
John does 2 jobs in 1 hour
jule does 4/3 job in 1 hour
Tom does 2/3 job in 1 hour

All theree will do 2+4/3 +2/3 of the job in 1 hour
4 jobs in 1 hour
for 1 job they will tale 1/4 hr or 15 minutes

Answer by richwmiller(17219)

x/30+x/45+x/90=1
3x+2x+x=90
6x=90
2x=30
x=15 minutes working together
check
15/30+15/45+15/90=1
1/2+1/3+1/6=1
3/6+2/6+1/6=1
John does 1/2 of the work three times Tom's portion
Julie does twice as much as Tom